OpenStudy (4meisu):

Factorize x^2 - 2xz - 2xy + 4yz

OpenStudy (whpalmer4):

Factor by grouping terms and factoring separately: \[(x^2-2xz) - (2xy - 4yz)\]\[x(x-2z) - 2y(x -2z)\] Can you factor that? Think distributive property in reverse...

OpenStudy (4meisu):

So I would have to factorize x(x-2z)?

OpenStudy (whpalmer4):

No, think of (x-2z) as a single item: We'll write \(a = x-2z\) Now our expression is \[xa - 2ya\] Can you factor that?

OpenStudy (4meisu):

a(x - 2y)

OpenStudy (whpalmer4):

Right! Now plug in the value of a and you have...

OpenStudy (4meisu):

I'm sorry, I got confused. Plug in the value of a with what? (Yes, I am super bad at Math)

OpenStudy (whpalmer4):

We simplified the recognition that you could factor \[x(x-2z) - 2y(x-2z)\] by writing \(a = x-2z\) and then factoring it. Now we need to replace a with \(x-2z\): \[a(x-2y) = (x-2z)(x-2y)\]

OpenStudy (whpalmer4):

You're not used to factoring out something like (x-2z), you're used to factoring out things like x, or 4. It takes a bit of practice to start spotting the more complicated factors.

OpenStudy (whpalmer4):

This procedure I went through is called factoring by grouping.

OpenStudy (4meisu):

Aaaah okay. Thank you!